# Microphysical Regulation of Non-Ideal MHD in Weakly-Ionized Systems: Does the Hall Effect Matter?

### v1

@@ 2.4.2 Generalization To Arbitrary Numbers of Species
Figure 3: Critical magnetic gradient scale-length $\ell_{\rm crit}$, below which the drift ($\ell_{\rm crit}^{\rm drift}$) or slip ($\ell_{\rm crit}^{\rm slip}$) speeds become superthermal, versus density $n$ for the same three model variants as Fig. 2. Strong gradients below this scale will be rapidly erased by enhanced resistivity (Fig. 2; § 4). For comparison we plot the disk scale-height or thermal-pressure scale-length $h_{T}$ predicted for a MMSN at the radii corresponding to each $n$. At circumstellar disk radii from $\sim 0.1-100\,$au (densities $n\sim 10^{10}-10^{18}{\rm cm^{-3}}$), and at all densities where the Hall diffusion would naively dominate absent anomalous correction terms (for $v_{\rm drift}\ll v_{T}$), we have $\ell_{\rm crit}\gg h_{T}$, so magnetic structures and Hall effects on disk scales will be strongly suppressed.
Given the full set of restricted assumptions in § 2.2, one can derive coefficients $\eta^{0}_{O,H,A}$ for an arbitrary number of species. In principle, as outlined in more detail in Appendix A, one could solve an entire chemical network for species abundances and simultaneously obtain the effective coefficients $\eta_{O,H,A}$ for an arbitrary set of velocity-dependent collision rates and anomalous resistivities $\omega^{\rm an}_{ij}$, based on the relative velocities between all pairs of species. But this would involve fundamental changes to the sorts of chemical networks and solvers usually employed, as well as introducing a number of ambiguities. Instead, much of the attention in modeling weakly ionized systems has focused on using chemical networks to calculate different charged-particle abundances $n_{i}$, which are then used in a fully operator-split manner to calculate $\eta^{0}_{O,H,A}$ from simplified expressions that depend only on $n_{i}$ (and background properties like $T$ or $B$, but not drift speeds). So it is important to consider how to generalize our approach to include many species. In this section, we develop a prescription that captures anomalous resistivity effects, but is sufficiently simple to be straightforwardly implemented into standard chemical-network methods.
@@
(Cowling, 1976; Nakano & Umebayashi, 1986; Wardle & Ng, 1999). The contribution of each to the current is then $|\delta{\bf J}|\sim n_{j}q_{j}\tilde{\beta}_{j}$. [^3] Note since this only appears as a weight function, the choice to multiply by $eB/c\rho$ is arbitrary. One can generalize Eq. 17 further for the case with arbitrary neutral density, where for the fully-ionized case the weights $\tilde{\beta}_{j}\propto q_{j}/m_{j}$, but in this case the various classical $\eta$ terms are generally negligible compared to the ideal MHD terms. Consider e.g. the simple case where there are two negative charge carriers – one “light” and relatively weakly collisionally coupled, and one “heavy” and much more strongly collisionally coupled (both with $q\sim-e$). This is often the case with electrons and charged grains, where the assumption of such a mass hierarchy is well justified. This gives the intuitive result that even if $n_{\rm grains}>n_{e}$, the current is still primarily carried by the electrons, unless $n_{\rm grains}/n_{e}\gg(m_{\rm grain}\gamma_{\rm grain})/(m_{e}\gamma_{en})\sim 10^{9.3}(R_{\rm grain}/{\rm\mu m})^{2}$. For the systems of interest (e.g. protostellar cores, circumstellar/protostellar disks, planetary atmospheres), the latter condition is rarely met, so even though grains can dominate the total charge, they do not actually carry most of the current. This means the current is being carried by relatively fewer free electrons, implying even higher drift speeds. [^4] Consider, for example, the limit of extremely large grains or “rocks.” Even if these somehow accumulated free electrons, clearly their electromotive forces are negligible and they would remain anchored to their positions, with Coulomb forces accumulating an opposite charge “locked” to their location. Thus even if they do not recombine, it is more appropriate to consider these effectively part of the neutral background, while the current and MHD dynamics of the fluid on scales large compared to the inter-grain separation are carried entirely by the remaining free electrons.
To calculate ${\bf v}_{\rm drift}$, we therefore take Eq. 9 (${\bf v}_{\rm drift}=(c\nabla\times\bar{\bf B})/(4\pi\bar{n}_{+}q_{+})=(c\nabla\times\bar{\bf B})/(4\pi\bar{n}_{-}|q_{-}|)$) and replace it with the weighted version:
@@
while $v_{\rm slip,\,\bot}$ (and hence $v_{\rm drift,\,max}$) can be calculated directly from Eq. 9 as $v_{\rm slip,\,\bot}=(\bar{\rho}/\bar{\rho}_{n})\,(\eta_{A}^{0}/\bar{B}^{2})\nabla\times\bar{\bf B}\times\bar{\bf B}$ using the total density and total density of (all) neutrals. Note that Eq. 18 immediately reduces to our expression from § 2.2 $\langle\bar{n}_{-}|q_{-}|\rangle=n_{-}|q_{-}|=n_{+}q_{+}$ when we just have two dominant charge carriers (one positive one negative), as it should. We similarly replace $n_{-}|q_{-}|$ or $n_{i}q_{i}$ by $\langle\bar{n}_{-}|q_{-}|\rangle$ (Eq. 18) in definitions like Eq. 13 for $\ell_{\rm crit}$.
For $\eta_{O}^{\rm an}$, note that we can now write:
$\eta_{O}^{\rm an}\sim\frac{c\bar{B}}{4\pi\langle\bar{n}_{-}|q_{-}|\rangle}\left[1+\frac{4\pi\bar{\rho}_{-}c^{2}}{\bar{B}^{2}}\right]^{1/2}\ ,$ (19)
so simply make the same replacement for $\bar{n}_{-}q_{-}\rightarrow\langle\bar{n}_{-}|q_{-}|\rangle\sim en_{e}$ in the prefactor, and note $\bar{\rho}_{-}\equiv m_{-}\langle n_{-}\rangle\sim m_{e}n_{e}$ should reflect the same dominant current-carrying species as $\langle\bar{n}_{-}|q_{-}|\rangle$. Also note that in the multi-species treatment, the pre-factors $\propto(n_{-}|q_{-}|)^{-1}$ in Eq. 15 are replaced by a similar weighted factor, so the two are consistent.
The generalization of $v_{T}$ is less obvious. Of course per § 2.3 the exact correct value here depends on the exact parameters and microphysics by way of exactly which instabilities and collision and interaction terms dominate on which scales. We can define an effective isothermal soundspeed for a multi-component plasma as:
@@
Naively, a sensible weight might be $\psi_{j}\sim$ constant for charged species and $=0$ for neutrals, or $\psi_{j}=|q_{j}|$. For an ion-electron plasma, this would give the expected answer of the thermal-ion speed (up to an $\mathcal{O}(1)$ prefactor, which is degenerate with our uncertainties in the “most salient” $v_{T}$ in any case). But this choice of $\psi$ would not correctly capture the case where one or both carriers are dust grains: with that definition $v_{T}\sim(k_{B}T/m_{\rm grain})^{1/2}\sim 0.01\,{\rm cm\,s^{-1}}\,(T/10\,{\rm K})^{1/2}\,(a_{\rm grain}/{\rm\mu m})^{-3/2}$ would become vanishingly small. This produces a number of results that might technically be valid if the only species were dust grains (i.e. no neutrals or ions or electrons present – an extreme “dusty plasma” case) but do not make sense in the context of interest here (where the dust thermal speed is extremely slow compared to that of other important species in the plasma). We can formulate a $\psi$ to capture realistic dusty plasmas by noting that in the limits of interest where dust is a dominant charge carrier, we are always in the weakly-ionized limit (large neutral density) with the dust strongly coupled to said neutrals ($\omega^{\rm coll}_{{\rm dust},n}\gg\Omega_{{\rm dust}}$). This means the neutrals should be included in Eq. 20, so $v_{T}\sim v_{T,n}$, and instabilities are excited when $v_{\rm drift}$ exceeds the neutral thermal speed. This also accords with the analysis of dust streaming instabilities/RDIs in Squire & Hopkins (2018b, a); Hopkins & Squire (2018b), where $v_{T,n}$ is the speed grains must be moving to excite strong instabilties on small scales. We can capture these cases appropriately by taking:
$\psi_{j}$$\equiv\begin{cases}|q_{j}|&(q_{j}\neq 0)\\ \frac{1}{\sum_{i}|q_{i}|n_{i}}\sum_{i}\frac{\omega^{\rm coll}_{ij}|q_{i}|n_{i}}{\omega^{\rm coll}_{ij}+\Omega_{i}}&(q_{j}=0)\ .\end{cases}$ (21)
Note the latter sum is nearly identical to sums that already need to be computed for $\eta_{O,H,A}^{0}$ in the standard multi-species formalism (Wardle & Ng, 1999) so entails no additional computational expense. In the weakly-ionized limit with any modest neutral coupling, this will tend to $v_{T}\rightarrow v_{T,n}$, while in the strongly ionized limit, this will tend to the thermal ion speed, as desired.
@@ 2.4.3 Specific Examples for Calculation
The relative values of the standard resistivity ($\eta$) coefficients are shown in Figs. 1, 2, & 3, for relatively simple assumptions of a multi-species system. In order to calculate actual values of the coefficients, of course, we need to make a number of specific assumptions about the scalings of $B=|{\bf B}|$, $n\equiv\rho/m_{p}$, ionization fractions, species abundances, temperatures, etc. These figures are intended for illustrative purposes only, so we adopt an extremely simple toy model for the quantities needed to compute the values shown. Specifically, we take (with all units here in CGS) similar parameters to those usually assumed in models of initial “warm” protostellar core collapse with temperature $T=10\sqrt{1+(n/10^{n_{T,1}})^{0.8}}(1+n/10^{n_{T,2}})^{-0.3}(1+n/10^{n_{T,3}})^{0.57}$ for $(n_{T,1},n_{T,2},n_{T,3})=(11,16,21)$ from Machida et al. (2006); and $B=10^{-6}n^{1/2}$ for $n<n_{0,B}$, $B=10^{-6}(nn_{0,B})^{1/4}$ for $n\geq n_{0,B}$ with $n_{0,B}=10^{6}$ from Wardle & Ng (1999). We adopt a simple five-species model of: molecular neutrals with Solar abundances ($\xi_{n}\approx 1$; $m_{n}=2.3m_{p}$); electrons ($n_{e}\sim 0.01$ for $n<10^{n_{e,1}}$ and $n_{e}\sim 0.01(n/10^{n_{e,1}})^{2}$ for $n\geq 10^{n_{e,1}}$, with $n_{e,1}=14$); atomic+molecular ions ($n_{i}=n_{e}$, $q_{i}=e$, $\langle m_{i}\rangle=29m_{p}$); negative and positive small dust grains ($n_{g+}=n_{g-}\sim 10^{-13}n$, $q_{g+}=-q_{g-}=e$, $R_{\rm grain}=3\times 10^{-6}$). These fitting functions give a rough approximation to the results from the chemical model using similar temperature and $B$ assumptions from NICIL (Wurster et al., 2021). And we take the zero-drift cross sections to have values $\langle\sigma v\rangle_{in}=1.5\times 10^{-9}$, $\langle\sigma v\rangle_{en}=4.7\times 10^{-10}T^{1/2}$, $\langle\sigma v\rangle_{gn}=5.1\times 10^{-6}T^{1/2}$, $\langle\sigma v\rangle_{ei}=51T^{-3/2}$ (Spitzer & Härm, 1953) and neglect any other reactions.
In addition to the above, we compare with two alternative models. The first is identical except we assume $B$ continues to rise more rapidly with $n$ (setting $n_{0,B}=10^{22}$). For the second we adopt temperatures and ionization fractions more appropriate for a circumstellar disk, specifically adjusting the model parameters above to give similar results to the minimum mass Solar nebula (MMSN) model from Chiang & Youdin (2010), taking $(n_{T,1},n_{T,2},n_{T,3})\rightarrow(10.5,12.5,19.5)$, $n_{e,1}\rightarrow 20$, and $n_{g\pm}\rightarrow 10^{-15}n$ at $n<10^{n_{e,1}}$.
@@
In Fig. 1, we estimate which of the Ohmic/Hall/ambipolar coefficients is largest in magnitude as a function of $B$ and $n$ (taking just the default model for abundances, $T$, etc., above, but allowing $B$ to freely vary at fixed values of all other properties at each $n$). In Fig. 2 and Fig. 3, we show the values of $\eta_{O,H,A}$ and the critical gradient length scales $\ell_{\rm crit}^{\rm drift}$ and $\ell_{\rm crit}^{\rm slip}$ (Eq. 13) for each of the three models above. In Figs. 12, we compare the standard non-ideal formulation for $\eta_{O,H,A}=\eta_{O,H,A}^{0}$ (Eq. 8), which ignores anomalous corrections and is therefore only valid when $\ell_{B}\gg\ell_{\rm crit}$ ($v_{\rm drift}\ll v_{T}$), to the corrected coefficients for the case of an $\ell_{B}$ which would otherwise give $v_{\rm drift}\sim 10\,v_{T}$ (using the implied $v_{\rm drift}$ from the standard formulation, Eq. 9), i.e. $\ell_{B}\sim 0.1\ell_{\rm crit}^{\rm drift}$. So Fig. 3 plots the dividing line between these cases, in terms of $\ell_{B}$.

### v2

@@ 2.4.2 Generalization To Arbitrary Numbers of Species
Figure 3: Critical magnetic gradient scale-length $\ell_{\rm crit}$, below which the drift ($\ell_{\rm crit}^{\rm drift}$) or slip ($\ell_{\rm crit}^{\rm slip}$) speeds become superthermal, versus density $n$ for the same three model variants as Fig. 2. Strong gradients below this scale will be rapidly erased by enhanced resistivity (Fig. 2; § 4). For comparison we plot the disk scale-height or thermal-pressure scale-length $h_{T}$ at each $n$ in the models. At circumstellar disk radii from $\sim 0.1-100\,$au (densities $n\sim 10^{10}-10^{18}{\rm cm^{-3}}$), and at densities where the Hall diffusion would naively dominate absent anomalous correction terms (for $v_{\rm drift}\ll v_{T}$), we have $\ell_{\rm crit}\gg h_{T}$, so magnetic structures and Hall effects on disk scales will be strongly suppressed.
Given the full set of restricted assumptions in § 2.2, one can derive coefficients $\eta^{0}_{O,H,A}$ for an arbitrary number of species. In principle, as outlined in more detail in Appendix A, one could numerically solve an entire chemical network for species abundances and simultaneously obtain the effective coefficients $\eta_{O,H,A}$ for an arbitrary set of velocity-dependent collision rates and anomalous resistivities $\omega^{\rm an}_{ij}$, based on the relative velocities between all pairs of species. But this would involve fundamental changes to the sorts of chemical networks and solvers usually employed, as well as introducing a number of ambiguities. Instead, much of the attention in modeling weakly ionized systems has focused on using chemical networks to calculate different charged-particle abundances $n_{i}$, which are then used in a fully operator-split manner to calculate $\eta^{0}_{O,H,A}$ from simplified expressions that depend only on $n_{i}$ (and background properties like $T$ or $B$, but not drift speeds). So it is important to consider how to generalize our approach to include many species. In this section, we develop a prescription that captures anomalous resistivity effects with sufficiently simple closed-form expressions that can be straightforwardly applied to the outputs of standard chemical-network methods (i.e. using information already supplied by those codes, rather than modifying the chemical computation itself, as we discuss in Appendix A, where we validate these expressions against the numerical approach).
@@
(Cowling, 1976; Nakano & Umebayashi, 1986; Wardle & Ng, 1999). The contribution of each to the current is then $|\delta{\bf J}|\propto n_{j}q_{j}\tilde{\beta}_{j}$. [^3] Note since this only appears as a weight function, the choice to multiply by $eB/c\rho$ is arbitrary. One can generalize Eq. 17 further for the case with arbitrary neutral density, where for the fully-ionized case the weights $\tilde{\beta}_{j}\propto q_{j}/m_{j}$, but in this case the various classical $\eta$ terms are generally negligible compared to the ideal MHD terms. Consider e.g. the simple case where there are two negative charge carriers – one “light” and relatively weakly collisionally coupled, and one “heavy” and much more strongly collisionally coupled (both with $q\sim-e$). This is often the case with electrons and charged grains, where the assumption of such a mass hierarchy is well justified. This gives the intuitive result that even if $n^{-}_{\rm grains}>n_{e}$, the current is still primarily carried by the electrons, unless $n^{-}_{\rm grains}/n_{e}\gg(m_{\rm grain}\gamma_{\rm grain})/(m_{e}\gamma_{en})\sim 10^{9.3}(R_{\rm grain}/{\rm\mu m})^{2}$. For the systems of interest (e.g. protostellar cores, circumstellar/protostellar disks, planetary atmospheres), the latter condition is rarely met, so even though grains can dominate the total charge, they do not actually carry most of the current. This means the current is being carried by relatively fewer free electrons, implying even higher drift speeds. [^4] Consider, for example, the limit of extremely large grains or “rocks.” Even if these somehow accumulated free electrons, clearly their electromotive forces are negligible and they would remain anchored to their positions, with Coulomb forces accumulating an opposite charge “locked” to their location. Thus even if they do not recombine, it is more appropriate to consider these effectively part of the neutral background, while the current and MHD dynamics of the fluid on scales large compared to the inter-grain separation are carried entirely by the remaining free electrons.
In that limit, the “standard” Ohmic coefficient then scales as $\eta_{O}^{0}=(cB/4\pi)/(\sum n_{j}q_{j}\tilde{\beta}_{j})$, i.e. is determined by the current-carrying species (the species which contributes the largest $|\delta{\bf J}|$). To determine an effective ${\bf v}_{\rm drift}$ of the dominant current carriers then, in order to estimate an effective Ohmic correction, we do not need to consider the relative drift velocity of every possible pair of charged species (thousands or millions of possible combinations), but the same weighted average. We therefore take Eq. 9 (${\bf v}_{\rm drift}=(c\nabla\times\bar{\bf B})/(4\pi\bar{n}_{+}q_{+})=(c\nabla\times\bar{\bf B})/(4\pi\bar{n}_{-}|q_{-}|)$) and replace it with the weighted version:
@@
Meanwhile $v_{\rm slip,\,\bot}$ (and hence $v_{\rm drift,\,max}$) can be calculated directly from Eq. 9 as ${\bf v}_{\rm slip,\,\bot}=(\bar{\rho}/\bar{\rho}_{n})\,(\eta_{A}^{0}/\bar{B}^{2})\nabla\times\bar{\bf B}\times\bar{\bf B}$ (unchanged), using the total density and total density of (all) neutrals. Since this builds in the weighting for multiple species already present in $\eta_{O}$ (defined in § A), this again ensures the proper weight (here towards slip velocity between the neutrals and the important inertia-bearing charged species). Note that Eq. 18 immediately reduces to our expression from § 2.2 $\langle\bar{n}_{-}|q_{-}|\rangle=n_{-}|q_{-}|=n_{+}q_{+}$ when we just have two dominant charge carriers (one positive one negative), as it should. We similarly replace $n_{-}|q_{-}|$ or $n_{i}q_{i}$ by $\langle\bar{n}_{-}|q_{-}|\rangle$ (Eq. 18) in definitions like Eq. 13 for $\ell_{\rm crit}$.
For $\eta_{O}^{\rm an}$, the salient anomalous resistivity is that of the same species that dominates the weighted average in our drift/current/Ohmic expression, so we can write:
$\eta_{O}^{\rm an}\sim\frac{c\bar{B}}{4\pi\langle\bar{n}_{-}|q_{-}|\rangle}\left[1+\frac{4\pi\langle\bar{\rho}_{-}\rangle c^{2}}{\bar{B}^{2}}\right]^{1/2}\ .$ (19)
This simply makes the same replacement for $\bar{n}_{-}q_{-}\rightarrow\langle\bar{n}_{-}|q_{-}|\rangle\sim en_{e}$ in the prefactor, and defines $\bar{\rho}_{-}\equiv m_{-}\langle n_{-}\rangle\sim m_{e}n_{e}$ to reflect the same dominant current-carrying species as $\langle\bar{n}_{-}|q_{-}|\rangle$. Also note that in the multi-species treatment, the pre-factors $\propto(n_{-}|q_{-}|)^{-1}$ in Eq. 15 are replaced by a similar weighted factor, so the two are consistent.
The generalization of $v_{T}$ as a “threshold” for anomalous behavior for the multi-species case is slightly less obvious. Of course, per § 2.3 the exact correct value here depends on the exact parameters and microphysics by way of exactly which instabilities and collision and interaction terms dominate on which scales. We can generically define an effective isothermal soundspeed for a multi-component plasma as:
@@
Naively, a sensible weight might be $\psi_{j}\sim$ constant for charged species and =0=0 for neutrals, or $\psi_{j}=|Z_{j}|=|q_{j}/e|$. For an ion-electron plasma, this would give the expected answer with $v_{T}$ the thermal-ion speed (up to an $\mathcal{O}(1)$ prefactor, which is degenerate with our uncertainties in the “most salient” $v_{T}$ in any case). But this choice of $\psi$ would not correctly capture the case where one or both carriers are dust grains: with that definition $v_{T}\sim(k_{B}T/m_{\rm grain})^{1/2}\sim 0.01\,{\rm cm\,s^{-1}}\,(T/10\,{\rm K})^{1/2}\,(a_{\rm grain}/{\rm\mu m})^{-3/2}$ would become vanishingly small. This produces a number of results that might technically be valid if the only species were dust grains (i.e. no neutrals or ions or electrons present – an extreme “dusty plasma” case) but do not make sense in the context of interest here (where the dust thermal speed is extremely slow compared to that of other important species in the plasma). We can formulate a $\psi$ to capture realistic dusty plasmas by noting that in the limits of interest where dust is a dominant charge carrier, we are always in the weakly-ionized limit (large neutral density) with the dust strongly coupled to said neutrals ($\omega^{\rm coll}_{{\rm dust},n}\gg\Omega_{{\rm dust}}$). This means the neutrals should be included in Eq. 20 since their inertia modifies the effective wavespeed of perturbations to the dust, so $v_{T}\sim v_{T,n}$, and instabilities are excited when $v_{\rm drift}$ exceeds the neutral thermal speed. This also accords with the formal analysis of dust streaming instabilities/RDIs in Squire & Hopkins (2018b, a); Hopkins & Squire (2018b), where $v_{T,n}$ is the speed grains must be moving to excite strong instabilties on small scales. We can capture these cases approximately by taking:
$\psi_{j}$ $\equiv\begin{cases}|q_{j}/e|&(q_{j}\neq 0)\\ \frac{1}{\sum_{i}|q_{i}|n_{i}}\sum_{i}\frac{\omega^{\rm coll}_{ij}|q_{i}|n_{i}}{\omega^{\rm coll}_{ij}+\Omega_{i}}&(q_{j}=0)\ .\end{cases}$ (21)
Note the latter sum is nearly identical to sums that already need to be computed for $\eta_{O,H,A}^{0}$ in the standard multi-species formalism (Wardle & Ng, 1999) so entails no additional computational expense. This captures the intuitive behavior in the limits of interest (and, as we show in § A, the numerical behavior in more exact computations): in the weakly-ionized limit with any modest neutral coupling, this will tend to $v_{T}\rightarrow v_{T,n}$, while in the strongly ionized limit, this will tend to the thermal ion speed, as desired.
@@ 2.4.3 Specific Examples for Calculation
In Figs. 1, 2, & 3, we illustrate the scalings derived above, showing relative values of $\eta_{O,\,H,\,A}$ and $\ell_{\rm crit}$ in some simple examples. In order to calculate specific values, we require a specific model for the chemistry (relative abundances of different charged species), density, temperature, magnetic field strength, etc. To do so, we consider a few different “default” models taken from the NICIL code (Wurster et al., 2021), a widely-used chemical network designed to compute the coefficients $\eta_{O,\,H,\,A}^{0}$ in mostly-neutral gas, with a large number of different ionized species, as well as both positively and negatively charged grains obeying an assumed grain size distribution. The coefficients in NICIL are calculated using the standard assumptions in § 2.2, notably assuming negligible drift, so we compare them to the corrected coefficients and scalings with our expressions, using the values of $T$, $B$, $n_{j}$, etc. output by the code. Specifically we consider (1) their “non-ideal MHD protostellar collapse” model and (2) their “baryotropic equation-of-state collapse” model (Wurster & Bate, 2019), and (3) a “protoplanetary disk” model (with a Mathis et al. 1977-like grain size distribution, attenuated cosmic ray ionization rate, and disk extended from $\sim 0.01-300$ au; Wurster et al. 2021). These make different assumptions for the scalings of $T$, $B$, and ionization fractions with density so give different coefficients relevant in different regimes. We stress that we are not arguing for any particular chemical model or scenario among these, but simply using them to illustrate the qualitative, systematic effects of superthermal drift in different situations. To this end, we have also re-computed Figs. 2-3 using a much more simplified five-species chemical model described in Appendix A. That gives different absolute values of the coefficients $\eta_{O,\,H,\,A}$, but the qualitative conclusions and systematic effects of the corrections from superthermal drift are the same in all cases.
As noted, NICIL considers a large number of species, with e.g. charged dust grains and many different ions relevant in dense gas. Thus our simple three-species model alone in § 2.1 does not un-ambiguously define “which” values of $v_{T}$ and other salient quantities to use in calculating the correction terms. For these, we specifically follow the generalized multi-species equations in § 2.4.2 above (Eqs. 19-20 for the effective $\eta_{O}^{\rm an}$ and $v_{T}$, respectively), with the dust included alongside the gas using the same grain sizes assumed in the NICIL calculation.
@@
In Fig. 1, we estimate which of the Ohmic/Hall/ambipolar coefficients is largest in magnitude as a function of $B$ and $n$ (taking just a simple model for $T(n)$ from Machida et al. (2006) and allowing $B$ to freely vary at fixed values of all other properties at each $n$, as our inputs to NICIL). In Fig. 2 and Fig. 3, we show the values of $\eta_{O,H,A}$ and the critical gradient length scales $\ell_{\rm crit}^{\rm drift}$ and $\ell_{\rm crit}^{\rm slip}$ (Eq. 13) for each of the three models above. In Figs. 1 & 2, we compare the standard non-ideal formulation for $\eta_{O,H,A}=\eta_{O,H,A}^{0}$ (Eq. 8), which ignores anomalous corrections and is therefore only valid when $\ell_{B}\gg\ell_{\rm crit}$ ($v_{\rm drift}\ll v_{T}$), to the corrected coefficients for the case of an $\ell_{B}$ which would otherwise give $v_{\rm drift}\sim 10\,v_{T}$ (using the implied $v_{\rm drift}$ from the standard formulation, Eqs. 9 & 18), i.e. $\ell_{B}\sim 0.1\ell_{\rm crit}^{\rm drift}$. So Fig. 3 plots the dividing line between these cases, in terms of $\ell_{B}$.